# Segregated Solutions to Critical Elliptic Systems in High Dimensions ($N\geq 5$)

### v1

@@ 1. Introduction
$\begin{cases}-\mathrm{i}\frac{\partial\Phi_{j}}{\partial t}=\Delta\Phi_{j}+\mu_{j}\lvert\Phi_{j}\rvert^{p-2}\Phi_{j}+\sum_{i\neq j}\beta_{ij}|\Phi_{i}|^{\frac{p}{2}}\lvert\Phi_{j}\rvert^{\frac{p}{2}-2}\Phi_{j},&~y\in\mathbb{R}^{N},~t>0,\\[8.53581pt] \Phi_{j}(y,0)=\phi_{j}(y),&~\ j=1,\dots,n.\end{cases}$ (1.2)
@@
To obtain the solitary wave solutions of system (1.1), one sets $\Phi_{j}(y,t)=e^{\mathrm{i}\lambda_{j}t}u_{j}(y)$, reducing the system to the to the following $n$-coupled nonlinear Schrödinger equations
@@
In recent years, significant mathematical studies have been conducted on nonlinear Schrödinger equations with critical exponents, for example, [2, 13, 22, 23, 24, 36, 37, 38, 41]. Del Pino, Musso, Pacard and Pistoia [15, 16] developed sequences of sign-changing solutions with high energy that concentrate along certain special sub-manifolds of $\mathbb{R}^{N}$. In [20], Guo, Li and Wei demonstrated the existence of infinitely many positive non-radial solutions to the coupled elliptic system. In [33], an unbounded sequence of non-radial positive vector solutions of segregated type was constructed when $p<\frac{2N}{N-2}$. In [34], Pistoia and Vaira found positive non-radial solutions for NN-coupled systems with weak interspecies forces in $N=2$ and $N=3$. For more information on the subcritical case of the system, we refer to [5, 13, 22, 24, 32, 38]. Recently, Chen, Medina and Pistoia [10] studied the existence of segregated solutions for a critical elliptic system with a small interspecies repulsive force in $\mathbb{R}^{4}$. When $n=1$ and $p=2^{*}$, system (1.3) reduces to the scalar equation
@@
Consequently, the system has infinitely many segregated solutions. Let us outline the main idea in the proof of Theorem $1.$ We additionally establish an alternative form of segregated solutions. In other words, segregated solutions display multiple bumps near infinity, with the bumps of $u_{1}$ and $u_{2}$ occurring on different circles. To illustrate this phenomenon, we first introduce some notations.
@@ 2. Preliminaries
$c_{1}\int\chi_{0}\left|\frac{\partial U_{0,\lambda}}{\partial y_{1}}\right|^{2}=\Big\langle-\Delta\varphi-(2^{*}-1)K_{1}\Big(\frac{\lvert y\rvert}{\mu}\Big)U^{2^{*}-2}_{r,\lambda}\varphi,\frac{\partial U_{x^{1},\lambda}}{\partial y_{1}}\Big\rangle-\langle 1_{P_{2}}h_{1},\frac{\partial U_{x^{1},\lambda}}{\partial y_{1}}\rangle.$ (2.4)
@@
To obtain decay estimates, we need the following comparison principle in $\Omega\subset\mathbb{R}^{N}$. For this purpose, we first recall the classical Sobolev embedding theorem. For every $N\geq 3$, there exists an optimal constant $\mathcal{S}>0$ depending only on NN such that
@@
where $C>0$ depends only on NN. Taking limits as $R\to\infty$, we obtain $\mathcal{S}\|\omega\|_{L^{2^{*}}(\Omega)}^{2}\leq\|\phi\rVert_{L^{\frac{N}{2}}(\Omega)}\|\omega\|^{2}_{L^{2^{*}}(\Omega)}$. Since $\|\phi\|_{L^{N/2}(\Omega)}<\mathcal{S}$, this implies $\|\omega\|_{L^{2^{*}}(\Omega)}=0$, so $\omega=0$ in $\Omega$. Thus, $v_{2}\leq v_{1}$ in $\Omega$. ∎
@@ 3. Minimization Problem
$\Lambda_{k}=\Big\{(\varphi,\psi)\in\mathbb{E}\mid\lVert(\varphi,\psi)\rVert_{\infty}\leq\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma^{N-2}_{k}}\Big\},$
@@ 3.1. Decay Estimates for Solutions and Deviations from the Limit Profile
$\lvert u_{1}(y)\rvert=\lvert\varphi_{0}+U_{r,\lambda}\rvert\leq U_{r,\lambda}+\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma^{N-2}_{k}}\leq C\varpi_{r,N-2},\quad y\in\partial P_{1}.$ (3.5)
@@
For $y\in\mathbb{R}^{N}\setminus B_{\frac{r+\rho}{2}}(0)$, we have $|y-x^{j}|\geq\frac{\rho-r}{2}=\frac{|x^{j}-y^{j}|}{2}$. Hence
@@
In $P_{1}$, if we assume further that $\theta<\frac{(N-4)\tau_{1}}{2(N-2)}$, then there holds
$\lvert\varphi\rvert\leq\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma^{N-2}_{k}}=\frac{1}{\mu^{\frac{m}{2}+\frac{(N-4)\tau_{1}}{2(N-2)}}}\frac{1}{\sigma_{k}^{\frac{N}{2}}}=o\left(\frac{1}{\mu^{\frac{m}{2}+\theta}\sigma_{k}^{\frac{N}{2}}}\right)<M_{0}\omega.$
@@ 1. If $1-\tau_{1}<\frac{(N-4)(N-2-\tau_{1})}{(N-2)^{2}}$, we can prove
$\lvert w(y)\rvert\leq$ $C\sigma^{N-2}_{k}(\|w_{0}\|_{L^{\infty}(P_{1})}+\|v_{0}\|_{L^{\infty}(Q_{1})})\sum_{i=1}^{k}\left(\frac{1}{(1+\lvert y-x^{i}\rvert)^{N-2}}+\frac{1}{(1+\lvert y-y^{i}\rvert)^{N-2}}\right).$ (3.14)
@@
We can obtain the estimate for $\nabla v$ in a similar way. ∎
@@ 3.4. The Existence of Minimizer
Define a functional $\Gamma:\mathbb{S}(\varphi_{0},\psi_{0})\rightarrow\mathbb{R}$ as
@@
The solution to (3.1) can be obtianed by solving the following minimizing problem
@@ 4. Reduction
$\begin{cases}-\Delta\varphi-(2^{*}-1)K_{1}\big(\frac{\lvert y\rvert}{\mu}\big)U^{2^{*}-2}_{r,\lambda}\varphi=l_{1}+R_{1}(\varphi)+D_{1}(\varphi,\psi)+\sum_{j=1}^{2}c_{j}Y_{j},\\[8.53581pt] -\Delta\psi-(2^{*}-1)K_{2}\big(\frac{\lvert y\rvert}{\mu}\big)V^{2^{*}-2}_{\rho,\nu}\psi=l_{2}+R_{2}(\psi)+D_{2}(\varphi,\psi)+\sum_{j=1}^{2}d_{j}Z_{j},\\[8.53581pt] (\varphi,\psi)\in\mathbb{E}.\end{cases}$ (4.1)
@@
$\begin{cases}\chi_{1}=1,~\text{in}~\mathbb{R}^{N}\setminus\cup^{k}_{i=1}B_{\frac{3}{4}\sigma^{2}_{k}}(x^{i}),\\[8.53581pt] \chi_{1}=0,~\text{in}~\cup^{k}_{i=1}B_{\frac{1}{4}\sigma^{2}_{k}}(x^{i}),\\[8.53581pt] \lvert\nabla\chi_{1}\rvert\leq\frac{4}{\sigma^{2}_{k}},~\lvert\Delta\chi_{1}\rvert\leq\frac{8N}{\sigma^{4}_{k}}.\end{cases}~\text{and}~\begin{cases}\chi_{2}=1,~\text{in}~\mathbb{R}^{N}\setminus\cup^{k}_{i=1}B_{\frac{3}{4}\sigma^{2}_{k}}(y^{i}),\\[8.53581pt] \chi_{2}=0,~\text{in}~\cup^{k}_{i=1}B_{\frac{1}{4}\sigma^{2}_{k}}(y^{i}),\\[8.53581pt] \lvert\nabla\chi_{2}\rvert\leq\frac{4}{\sigma^{2}_{k}},~\lvert\Delta\chi_{2}\rvert\leq\frac{8N}{\sigma^{4}_{k}}.\end{cases}$ (4.2)
@@
$\leq$ $\sigma^{4}_{k}\lVert\varphi_{0}\Delta\chi_{1}+2\nabla\chi_{1}\nabla\varphi_{0}\rVert_{L^{\infty}(P_{2})}+\sigma^{4}_{k}\lVert l_{1}+R_{1}(\varphi_{0})\rVert_{L^{\infty}(P_{2})}\leq\frac{1}{2}\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma^{N-2}_{k}}.$
@@
$\lVert L_{2}\big(\psi_{0}\Delta\chi_{2}+2\nabla\chi_{2}\nabla\psi_{0}+(1-\chi_{2})\big(l_{2}+R_{2}(\psi_{0})\big)\big)\rVert_{L^{\infty}(\mathbb{R}^{N})}\leq\frac{1}{2}\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma^{N-2}_{k}}.$
@@
$\leq$ $\frac{C}{k\mu^{\frac{m}{2}}}(\lVert w_{0}\rVert_{L^{\infty}(P_{1})}+\|v_{0}\|_{L^{\infty}(Q_{1})}).$
@@
$\lVert R_{1}(\varphi_{0})-R_{1}(\bar{\varphi}_{0})\rVert_{L^{\infty}(P_{2})}\leq C\frac{1}{\mu^{\frac{2m}{N-2}}\sigma^{4}_{k}}(\lVert w_{0}\rVert_{L^{\infty}(P_{1})}+\|v_{0}\|_{L^{\infty}(Q_{1})}).$ (4.11)
@@
By Proposition 2.2 and Remark 3.11, we know $(\varphi_{0},\psi_{0})\in(D^{1,2}(\mathbb{R}^{N}))^{2}$ and $(\zeta_{1},\zeta_{2})\in(D^{1,2}(\mathbb{R}^{N}))^{2}$. Recalling the cut off functions $(\chi_{1},\chi_{2})$ defined in (4.2), and the definition of $S$, we obtain
@@
Hence $\zeta_{1}=\varphi_{0}$. Similarly, $\zeta_{2}=\psi_{0}$.
@@ 5. Proof of Main Theorems
Therefore, if the left hand side of (5.1) belongs to $\mathbb{E}$, then the function on the right hand side of (5.1) must be zero. Recall that
@@
$M_{4}$ $=-(2^{*}-1)\int_{\mathbb{R}^{N}}\left(K_{1}\left(\frac{\lvert y\rvert}{\mu}\right)-1\right)U_{x^{1},\lambda}^{2^{*}-2}\varphi\zeta\frac{\partial U_{x^{1},\lambda}}{\partial\lambda}.$
@@
$\leq$ $\frac{C}{\mu^{\frac{4m}{2N-3}+\frac{m}{2}+\theta}}\int_{B_{2\mu^{\frac{2m}{2N-3}}}(x^{1})\setminus B_{\mu^{\frac{2m}{2N-3}}}(x^{1})}\frac{1}{(1+|y-x^{1}|)^{\frac{3N}{2}-2}}\leq\frac{C}{\mu^{\frac{Nm}{2N-3}+\frac{m}{2}+\theta}}=O\Big(\frac{1}{\mu^{m+\theta}}\Big).$
@@
Proof of Theorem 1.2 and Theorem $1.$
@@
$\lambda_{0}=\Big(\frac{B_{5}}{B_{1}r^{N-2}_{0}}\Big)^{\frac{1}{N-2-m}},\quad\nu_{0}=\Big(\frac{B_{6}}{B_{3}\rho^{N-2}_{0}}\Big)^{\frac{1}{N-2-m}}.$
by the solution of
@@
$\mathbb{M}:=\Big[r_{0}\mu-\displaystyle\frac{1}{\mu^{\bar{\theta}}},r_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}\Big]\times\Big[\rho_{0}\mu-\displaystyle\frac{1}{\mu^{\bar{\theta}}},\rho_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}\Big]\times\Big[\lambda_{0}-\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}},\lambda_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}\Big]\times\Big[\nu_{0}-\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}},\nu_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}\Big].$
@@
Similarly, we can get $\widetilde{c}_{2}<0$ if $\lambda=\lambda_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}$; $\widetilde{d}_{2}>0$ if $\nu=\nu_{0}-\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}$; $\widetilde{d}_{2}<0$ if $\nu=\nu_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}$. $\widetilde{c}_{1}<0$ if $r=r_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}$; $\widetilde{d}_{1}>0$ if $\rho=\rho_{0}\mu-\displaystyle\frac{1}{\mu^{\bar{\theta}}}$ $\widetilde{d}_{1}<0$ if $\rho=\rho_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}$.

### v2

@@ 1. Introduction
$\begin{cases}-\mathrm{i}\frac{\partial\Phi_{j}}{\partial t}=\Delta\Phi_{j}+\mu_{j}\lvert\Phi_{j}\rvert^{p-2}\Phi_{j}+\sum_{i\neq j}\beta_{ij}|\Phi_{i}|^{\frac{p}{2}}\lvert\Phi_{j}\rvert^{\frac{p}{2}-2}\Phi_{j},&~y\in\mathbb{R}^{N},~t>0,\\[8.53581pt] \Phi_{j}(y,0)=\phi_{j}(y),&~\ j=1,\dots,n,\end{cases}$ (1.2)
@@
To obtain the solitary wave solutions of system (1.1), one sets $\Phi_{j}(y,t)=e^{\mathrm{i}\lambda_{j}t}u_{j}(y)$, reducing the system to the following $n$-coupled nonlinear Schrödinger equations
@@
In recent years, significant mathematical studies have been conducted on nonlinear Schrödinger equations with critical exponents, for example, [2, 13, 22, 23, 24, 36, 37, 38, 41]. Del Pino, Musso, Pacard and Pistoia [15, 16] developed sequences of sign-changing solutions with high energy that concentrate along certain special sub-manifolds of $\mathbb{R}^{N}$. In [20], Guo, Li and Wei demonstrated the existence of infinitely many positive non-radial solutions to the coupled elliptic system. In [33], an unbounded sequence of non-radial positive vector solutions of segregated type was constructed when $p<\frac{2N}{N-2}$. In [34], Pistoia and Vaira found positive non-radial solutions for $N$-coupled systems with weak interspecies forces in $N=2$ and $N=3$. For more information on the subcritical case of the system, we refer to [5, 13, 22, 24, 32, 38]. Recently, Chen, Medina and Pistoia [10] studied the existence of segregated solutions for a critical elliptic system with a small interspecies repulsive force in $\mathbb{R}^{4}$. When $n=1$ and $p=2^{*}$, system (1.3) reduces to the scalar equation
@@
Consequently, the system has infinitely many segregated solutions. Let us outline the main idea in the proof of Theorem 1.1. We additionally establish an alternative form of segregated solutions. In other words, segregated solutions display multiple bumps near infinity, with the bumps of $u_{1}$ and $u_{2}$ occurring on different circles. To illustrate this phenomenon, we first introduce some notations.
@@ 2. Preliminaries
$c_{1}(\int\chi_{0}\left|\frac{\partial U_{0,\lambda}}{\partial y_{1}}\right|^{2}+o(1))=\Big\langle-\Delta\varphi-(2^{*}-1)K_{1}\Big(\frac{\lvert y\rvert}{\mu}\Big)U^{2^{*}-2}_{r,\lambda}\varphi,\frac{\partial U_{x^{1},\lambda}}{\partial y_{1}}\Big\rangle-\langle 1_{P_{2}}h_{1},\frac{\partial U_{x^{1},\lambda}}{\partial y_{1}}\rangle.$ (2.4)
@@
To obtain decay estimates, we need the following comparison principle in $\Omega\subset\mathbb{R}^{N}$. For this purpose, we first recall the classical Sobolev embedding theorem. For every $N\geq 3$, there exists an optimal constant $\mathcal{S}>0$ depending only on $N$ such that
@@
where $C>0$ depends only on $N$. Taking limits as $R\to\infty$, we obtain $\mathcal{S}\|\omega\|_{L^{2^{*}}(\Omega)}^{2}\leq\|\phi\rVert_{L^{\frac{N}{2}}(\Omega)}\|\omega\|^{2}_{L^{2^{*}}(\Omega)}$. Since $\|\phi\|_{L^{N/2}(\Omega)}<\mathcal{S}$, this implies $\|\omega\|_{L^{2^{*}}(\Omega)}=0$, so $\omega=0$ in $\Omega$. Thus, $v_{2}\leq v_{1}$ in $\Omega$. ∎
@@ 3. Minimization Problem
$\Lambda_{k}=\Big\{(\varphi,\psi)\in\mathbb{E}\mid\lVert(\varphi,\psi)\rVert_{\infty}\leq\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma_{k}^{\frac{N+1}{2}}}\Big\},$
@@ 3.1. Decay Estimates for Solutions and Deviations from the Limit Profile
$\lvert u_{1}(y)\rvert=\lvert\varphi_{0}+U_{r,\lambda}\rvert\leq U_{r,\lambda}+\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma_{k}^{\frac{N+1}{2}}}\leq C\varpi_{r,N-2},\quad y\in\partial P_{1}.$ (3.5)
@@
For $y\in\mathbb{R}^{N}\setminus B_{\frac{r+\rho}{2}}(0)$, we have $|y-x^{j}|\geq\frac{\rho-r}{2}=\frac{|x^{j}-y^{j}|}{2}$. Hence,
@@
In $P_{1}$, if we assume further that $\theta<\frac{N-2-m}{2(N-2)^{2}}$, then there holds
$\lvert\varphi\rvert\leq\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma_{k}^{\frac{N+1}{2}}}=\frac{1}{\mu^{\frac{m}{2}+\frac{N-2-m}{2(N-2)^{2}}}}\frac{1}{\sigma_{k}^{\frac{N}{2}}}=o\left(\frac{1}{\mu^{\frac{m}{2}+\theta}\sigma_{k}^{\frac{N}{2}}}\right)<M_{0}\omega.$
@@ 1. If $1-\tau_{1}<\frac{(N-4)(N-2-\tau_{1})}{(N-2)^{2}}$, we can prove
$\lvert w(y)\rvert\leq$ $C\sigma^{N-2}_{k}(\|w_{0}\|_{L^{\infty}(P_{1})}+\|v_{0}\|_{L^{\infty}(Q_{1})})\sum_{i=1}^{k}\left(\frac{1}{(1+\lvert y-x^{i}\rvert)^{N-2}}+\frac{1}{(1+\lvert y-y^{i}\rvert)^{N-2}}\right),$ (3.14)
@@
We can obtain the estimate for $\nabla v$ similarly. ∎
@@ 3.4. The Existence of Minimizer
Define a functional $\Gamma\colon\mathbb{S}(\varphi_{0},\psi_{0})\rightarrow\mathbb{R}$ as
@@
The solution to (3.1) can be obtained by solving the following minimizing problem
@@ 4. Reduction
$\begin{cases}-\Delta\varphi-(2^{*}-1)K_{1}\big(\frac{\lvert y\rvert}{\mu}\big)U^{2^{*}-2}_{r,\lambda}\varphi=l_{1}+R_{1}(\varphi)+D_{1}(\varphi,\psi)+\sum_{j=1}^{2}c_{j}Y_{j},\\[8.53581pt] -\Delta\psi-(2^{*}-1)K_{2}\big(\frac{\lvert y\rvert}{\mu}\big)V^{2^{*}-2}_{\rho,\nu}\psi=l_{2}+R_{2}(\psi)+D_{2}(\varphi,\psi)+\sum_{j=1}^{2}d_{j}Z_{j},\\[8.53581pt] (\varphi,\psi)\in\mathbb{E},\end{cases}$ (4.1)
@@
$\begin{cases}\chi_{1}=1,~\text{in}~\mathbb{R}^{N}\setminus\cup^{k}_{i=1}B_{\frac{3}{4}\sigma^{2}_{k}}(x^{i}),\\[8.53581pt] \chi_{1}=0,~\text{in}~\cup^{k}_{i=1}B_{\frac{1}{4}\sigma^{2}_{k}}(x^{i}),\\[8.53581pt] \lvert\nabla\chi_{1}\rvert\leq\frac{4}{\sigma^{2}_{k}},~\lvert\Delta\chi_{1}\rvert\leq\frac{8N}{\sigma^{4}_{k}},\end{cases}~\text{and}~\begin{cases}\chi_{2}=1,~\text{in}~\mathbb{R}^{N}\setminus\cup^{k}_{i=1}B_{\frac{3}{4}\sigma^{2}_{k}}(y^{i}),\\[8.53581pt] \chi_{2}=0,~\text{in}~\cup^{k}_{i=1}B_{\frac{1}{4}\sigma^{2}_{k}}(y^{i}),\\[8.53581pt] \lvert\nabla\chi_{2}\rvert\leq\frac{4}{\sigma^{2}_{k}},~\lvert\Delta\chi_{2}\rvert\leq\frac{8N}{\sigma^{4}_{k}}.\end{cases}$ (4.2)
@@
$\leq$ $\sigma^{4}_{k}\lVert\varphi_{0}\Delta\chi_{1}+2\nabla\chi_{1}\nabla\varphi_{0}\rVert_{L^{\infty}(P_{2})}+\sigma^{4}_{k}\lVert l_{1}+R_{1}(\varphi_{0})\rVert_{L^{\infty}(P_{2})}\leq\frac{1}{2}\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma_{k}^{\frac{N+1}{2}}}.$
@@
$\lVert L_{2}\big(\psi_{0}\Delta\chi_{2}+2\nabla\chi_{2}\nabla\psi_{0}+(1-\chi_{2})\big(l_{2}+R_{2}(\psi_{0})\big)\big)\rVert_{L^{\infty}(\mathbb{R}^{N})}\leq\frac{1}{2}\frac{1}{\mu^{\frac{m}{2}}}\frac{1}{\sigma_{k}^{\frac{N+1}{2}}}.$
@@
$\leq$ $\frac{C}{\sigma_{k}^{\frac{N+1}{2}}\mu^{\frac{m}{2}}}(\lVert w_{0}\rVert_{L^{\infty}(P_{1})}+\|v_{0}\|_{L^{\infty}(Q_{1})}).$
@@
$\lVert R_{1}(\varphi_{0})-R_{1}(\bar{\varphi}_{0})\rVert_{L^{\infty}(P_{2})}\leq C\frac{1}{\mu^{\frac{2m}{N-2}}\sigma^{\frac{2(N+1)}{N-2}}_{k}}(\lVert w_{0}\rVert_{L^{\infty}(P_{1})}+\|v_{0}\|_{L^{\infty}(Q_{1})}).$ (4.11)
@@
By Proposition 2.2 and Remark 3.11, we know $(\varphi_{0},\psi_{0})\in(D^{1,2}(\mathbb{R}^{N}))^{2}$ and $(\zeta_{1},\zeta_{2})\in(D^{1,2}(\mathbb{R}^{N}))^{2}$. Recalling the cut-off functions $(\chi_{1},\chi_{2})$ defined in (4.2), and the definition of $S$, we obtain
@@
Hence, $\zeta_{1}=\varphi_{0}$. Similarly, $\zeta_{2}=\psi_{0}$.
@@ 5. Proof of Main Theorems
Therefore, if the left-hand side of (5.1) belongs to $\mathbb{E}$, then the function on the right-hand side of (5.1) must be zero. Recall that
@@
$M_{4}$ $=-(2^{*}-1)\int_{\mathbb{R}^{N}}\left(K_{1}\left(\frac{\lvert y\rvert}{\mu}\right)-1\right)U_{x^{1},\lambda}^{2^{*}-2}\varphi\zeta\frac{\partial U_{x^{1},\lambda}}{\partial\lambda},$
@@
$\leq$ $\frac{C}{\mu^{\frac{4m}{2N-3}+\frac{m}{2}+\theta}}\int_{B_{2\mu^{\frac{2m}{2N-3}}}(x^{1})\setminus B_{\mu^{\frac{2m}{2N-3}}}(x^{1})}\frac{1}{(1+|y-x^{1}|)^{\frac{3N}{2}-2}}\leq\frac{C}{\mu^{\frac{Nm}{2N-3}+\frac{m}{2}+\theta}}=O\Big(\frac{1}{\mu^{m+\theta}}\Big),$
@@
Proof of Theorem 1.2 and Theorem 1.1.
@@
$\lambda_{0}=\Big(\frac{B_{5}}{B_{1}r^{N-2}_{0}}\Big)^{\frac{1}{N-2-m}},\quad\nu_{0}=\Big(\frac{B_{6}}{B_{3}\rho^{N-2}_{0}}\Big)^{\frac{1}{N-2-m}}$
be the solution of
@@
$\mathbb{M}:=\Big[r_{0}\mu-\displaystyle\frac{1}{\mu^{\bar{\theta}}},r_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}\Big]\times\Big[\rho_{0}\mu-\displaystyle\frac{1}{\mu^{\bar{\theta}}},\rho_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}\Big]\times\Big[\lambda_{0}-\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}},\lambda_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}\Big]\times\Big[\nu_{0}-\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}},\nu_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}\Big],$
@@
Similarly, we can get $\widetilde{c}_{2}<0$ if $\lambda=\lambda_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}$; $\widetilde{d}_{2}>0$ if $\nu=\nu_{0}-\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}$; $\widetilde{d}_{2}<0$ if $\nu=\nu_{0}+\frac{1}{\mu^{\frac{3}{2}\bar{\theta}}}$. $\widetilde{c}_{1}<0$ if $r=r_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}$; $\widetilde{d}_{1}>0$ if $\rho=\rho_{0}\mu-\displaystyle\frac{1}{\mu^{\bar{\theta}}}$ $\widetilde{d}_{1}<0$; if $\rho=\rho_{0}\mu+\displaystyle\frac{1}{\mu^{\bar{\theta}}}$.